Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-037_2000-043

Mission Operations Days: 2000/037 to 2000/043
February 06, 2000 00:00:00 GMT to February 12, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA On (285 K); B OffNominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, ClosedNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA Off; B OnNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off, #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BDNominal
Inhibit Ids Set31,32,33,53,58BB, CE, CP, SDSM, SD
TMONs EnabledNoneNominal

This Week’s Completed MODIS Activities:

Sunday, February 6, 2000

None

Monday, February 7, 2000

None

Tuesday, February 8, 2000

None

Wednesday, February 9, 2000

None

Thursday, February 10, 2000

None

Friday, February 11, 2000

2000/042 15:44 Real time - - Transition to Standby A Mode from Safe Mode
2000/042 15:47 Real time - - SVD to Open from Closed
2000/042 15:51 Real time - - Transition to Science B Mode from Standby Mode
2000/042 16:34 Real time - - First science data dump – three uncorrectable EDU errors, expected only two. This was caused by starting with an empty SSR buffer.
2000/042 19:01:35 Real time - - PV Ecal run via real-time procedure, macro 25
2000/042 19:04:09 Real time - - PV Ecal macro 25 complete
2000/042 19:05 Real time - - BB on to 315K, B-side (setpoint = 141)
2000/042 19:07 Real time - - BB temp set to 285K (setpoint = 1703)
2000/042 20:45 Real time - - SDD opened to screened position, step = 1069
2000/042 20:50 Real time - - SDD closed, step = 0
2000/042 20:55 Real time - - Microprocessor dump of door table
2000/042 22:17:40 Real time - - SDD fully opened, step = 2151
2000/042 22:26:32 Real time - - SDD fully close, step = 0
2000/042 22:29 Real time - - Microprocessor dump of door table

Saturday, February 12, 2000

2000/043 12:26 - Replay of data was not done because buffer had already overwritten it. There will be a data gap here abouts.
2000/043 14:24 - Replay of data was not done because buffer had already overwritten it. There will be a data gap here abouts.

This Week's Scheduled MODIS Activities Not Completed:

2000/042 17:25 ATC - - PV Ecal did not occur because the inhibit ID for CP was still set.

Upcoming MODIS Events:

2000/043 20:41 - 150s orbit attainment burn
2000/044 16:48 Real time - - Turn off MODIS
Real time - - Turn on MODIS on A-side
Real time - - Transition to Science Mode on A-Side
2000/044 18:17 Real time - - SDSM functional test
2000/044 18:30 Real time - - Reset Inhibit Ids 31, 32, 33, 53, 58
2000/044 20:43:21 ATC - - OA13: EDS to support OA16 begins
2000/044 20:43:25 ATC - - OA16: SD/SDSM Screened start
2000/044 21:01:18 ATC - - OA16: SD/SDSM Screened end
2000/044 21:01:22 ATC - - OA13: EDS to support OA16 ends
2000/044 21:15:00 ATC - - OA28: PC Electronics Calibration start
2000/044 21:19:31 ATC - - OA28: PC Electronics Calibration end
2000/044 21:20:00 ATC - - OA27: PV Electronics Calibration start
2000/044 21:22:31 ATC - - OA27: PV Electronics Calibration end
2000/044 22:10 Real time - - Set BB duty cycle to 100%
2000/044 22:21:39 ATC - - OA13: EDS to support OA15 begins
2000/044 22:21:41 ATC - - OA15: SD/SDSM Open start
2000/044 22:39:36 ATC - - OA15: SD/SDSM Open end
2000/044 22:39:40 ATC - - OA13: EDS to support OA15 ends
2000/045 17:00 ATC - - OA26: BB On to 315K at 100% duty cycle
2000/045 20:12 - 300s orbit attainment burn
2000/046 01:00 ATC - - Blackbody turn off
2000/046 14:00 Real time - - Set BB duty cycle to 33%
2000/046 15:00 Real time - - SRCA 1W functional (PC17)
2000/046 17:00 ATC - OA26: BB to 315K at 33% duty cycle
2000/047 01:00 ATC - Blackbody turn off
2000/047 ATC – SD/SDSM Screened
2000/047 ATC – SD/SDSM Screened
2000/047 14:30:00 – 15:03:33 ATC - - SRCA Full Radiometric
2000/047 ATC – SD/SDSM Open
2000/047 19:46 - 300s orbit attainment burn
2000/048 14:43:00 – 15:40:31 ATC - - SRCA Full Spatial
2000/049 21:52 - 300s orbit attainment burn
2000/047 ATC – SD/SDSM Screened
2000/047 ATC – SD/SDSM Screened
2000/047 ATC - - SRCA Full Spectral
2000/047 ATC – SD/SDSM Open

MODIS Anomalies:

None.

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD open and BB heating to 285K. Solar Diffuser Door opened to the screened and full open position nominally. (And there was much rejoicing!)

Orbit attainment burns are proceeding nominally.

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

2000/041 21:15 - 60s orbit attainment burn
2000/043 12:26 - Replay of data was not done because buffer had already overwritten it. There will be a data gap here abouts.
2000/043 14:24 - Replay of data was not done because buffer had already overwritten it. There will be a data gap here abouts.

The following is the currently planned burn schedule to place Terra into her final orbit (the times are in GMT). Note that these times have been updated since yesterday’s report. Data loss is expected around every burn.

Burn | Start Time | End Time | Duration(sec)

1 | Feb 10 2000 21:15:43.030 | Feb 10 2000 21:16:43.030 | 60.000 (complete)
2 | Feb 12 2000 20:43:22.876 | Feb 12 2000 20:45:52.876 | 150.000
3 | Feb 14 2000 20:12:16.920 | Feb 14 2000 20:17:16.920 | 300.000
4 | Feb 16 2000 19:46:07.615 | Feb 16 2000 19:51:07.615 | 300.000
5 | Feb 18 2000 21:52:50.413 | Feb 18 2000 21:57:50.413 | 300.000
6 | Feb 20 2000 19:04:11.675 | Feb 20 2000 19:09:11.675 | 300.000
7 | Feb 22 2000 21:18:02.752 | Feb 22 2000 21:23:02.752 | 300.000
8 | Feb 24 2000 20:15:24.276 | Feb 24 2000 20:16:24.276 | 60.000

Limited Life Item Status:

All limited life items are well within lifetime ranges. The precise statistics for each item have been received from LMMS/Valley Forge and will be incorporated next week.